A. Solve for the unknown using the SCI format. Show your solution and answer.

Case 1:

Expenses P20,000

Net Income/(Net Loss) 50,000

Service Income ?

Case 2:

Net Income/(Net Loss) P67,500

Service Revenue 80,000

Expenses 12,500

Other Comprehensive Income 10,500

Comprehensive Income ?

Case 3:

Comprehensive Income P40,000

Other Comprehensive Income 4,000

Service Income 130,000

Net Income/(Net Loss) 36,000

Expenses ?


1
Expert's answer
2021-09-12T19:32:11-0400

Solution:

Case 1:

As per the SCI format:

Service Income less Expenses is equal to Net income:


Therefore,

Service income = Net income + Expenses = 50,000 + 20,000 = 70,000

Service income = 70,000


Service Income                    70,000

Less: Expenses                    (50,000)

Net Income                           20,000


Case 2:

Service Revenue                                 80,000

Less: Expenses                                   (12,500)

Net Income                                         67,500

Add: Other comprehensive income   10,500

Comprehensive Income                   78,000

 

Case 3:

Service Income                                   130,000

Less: Expenses                                   (94,000)

Net Income                                           36,000

Add: Other comprehensive income       4,000

Comprehensive Income                        40,000

Expenses = Service income – Net income = 130,000 – 36,000 = 94,000

Expenses = 94,000